OpenStudy (anonymous):

If a line crosses the y-axis at (0, -4) and has a slope of -2, what is the equation of the line?

OpenStudy (anonymous):

2x − y = -4 2x + y = -4 2x + y = 4 2x − y = 4

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

got it?

OpenStudy (igreen):

Hold on

OpenStudy (igreen):

B